When being unloaded from a moving truck, a 15.5- kilogram suitcase is placed on a flat ramp inclined at 33.0 o. When released from rest, the suitcase accelerates down the ramp at 1.32 m/s². What is the coefficient of kinetic friction between the suitcase and the ramp?
Answer #1

Here,

let the coefficient of kinetic friction is u

Now, using second law of motion

net force = mass * acceleration

15.5 * 1.32 = m * g * sin(theta) - u *m * g * cos(theta)

15.5 * 1.32 = 15.5 * 9.8 * sin(33 degree) - u * 15.5 * 9.8 * cos(33 degree)

solving for u

u = 0.49

the coefficient of kinetic friction is 0.49